OpenStudy (anonymous):

A particle is moving along the x-axis so that its position at t ≥ 0 is given by s(t) = (t)In(3t). Find the acceleration of the particle when the velocity is first zero.

OpenStudy (solomonzelman):

You first need to find the point where the velocity of the particle is zero, and then to find the acceleration at this point (when the velocity is 0).

OpenStudy (solomonzelman):

The position is: \(\color{#000000 }{ \displaystyle s(t)=t\ln (3t) }\)

OpenStudy (solomonzelman):

The slope of the position graph tells you how fast the function is going. Or, s'(t) would tell you how fast the particle is moving at each instant "t".

OpenStudy (solomonzelman):

So you need to find the derivative: \(\color{#000000 }{ \displaystyle s'(t)=}\) (Use the product rule)

OpenStudy (anonymous):

Ummm I don't know @solomonzelman

OpenStudy (solomonzelman):

Be specific please. What don't you know?

OpenStudy (anonymous):

Would the derivative be ln(3t)+1?

OpenStudy (anonymous):

@solomonzelman

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle s'(t)=t\times \left(\ln(3t)\right)'+\left(t\right)'\times \ln(3t)}\) \(\color{#000000 }{ \displaystyle s'(t)=t\times \left(\frac{1}{3t}\times 3\right)+1\times \ln(3t)}\) \(\color{#000000 }{ \displaystyle s'(t)=1+ \ln(3t)}\)

OpenStudy (solomonzelman):

Very good!

OpenStudy (solomonzelman):

s'(t) is the velocity, so you can say, v(t)=1+ln(3t)

OpenStudy (solomonzelman):

Now, when is the velocity of the particle 0? (at what values of t?)

OpenStudy (solomonzelman):

to answer this question set v(t)=0

OpenStudy (anonymous):

Ok thanks

OpenStudy (solomonzelman):

And that is not going to be the end.

OpenStudy (solomonzelman):

Can you solve 0=ln(3t)+1 ?

OpenStudy (anonymous):

I know you solve it thanks that's all the help I needed

OpenStudy (solomonzelman):

Just to make sure \( ...\) After you solve the above equation for t, you will need to find the acceleration at this point. The derivative of velocity is acceleration, so v'(t) , (or s''(t)). Find s''(t), and then plug in that value of t, at which s'(t) (or v(t)) =0.